OpenStudy (anonymous):

Simplify. square root of 3 times square root of 21 square root of 24 square root of 63 3 square root of 7 None of the above

OpenStudy (anonymous):

@lizz123

OpenStudy (lizz123):

http://gyazo.com/bd2a1f0a808cf7ac9a5d4dbd572b3bc2

Nnesha (nnesha):

\[\sqrt{3} \times \sqrt{21}\] multiply both numbers

Nnesha (nnesha):

3 times 21 = ??

OpenStudy (anonymous):

63 @Nnesha

Nnesha (nnesha):

yes but under square root \[\sqrt{63}\] now you have to factor this

OpenStudy (anonymous):

the answer is c???

Nnesha (nnesha):

square root of 9 is 3 so yes